# CryoSplat: Gaussian Splatting for Cryo-EM Homogeneous Reconstruction

### v2

@@ Abstract
As a critical modality for structural biology, cryogenic electron microscopy (cryo-EM) facilitates the determination of macromolecular structures at near-atomic resolution. The core computational task in single-particle cryo-EM is to reconstruct the 3D electrostatic potential of a molecule from a large collection of noisy 2D projections acquired at unknown orientations. Gaussian mixture models (GMMs) provide a continuous, compact, and physically interpretable representation for molecular density and have recently gained interest in cryo-EM reconstruction. However, existing methods rely on external consensus maps or atomic models for initialization, limiting their use in self-contained pipelines. Addressing this issue, we introduce cryoSplat, a GMM-based method that integrates Gaussian splatting with the physics of cryo-EM image formation. In particular, we develop an orthogonal projection-aware Gaussian splatting, with adaptations such as a normalization term and FFT-aligned coordinate system tailored for cryo-EM imaging. All these innovations enable stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images using random initialization. Experimental results on real datasets validate the effectiveness and robustness of cryoSplat over representative baselines. The code will be released upon publication.
footnotetext: Preprint. Under review.
Introduction
Single particle cryogenic electron microscopy (cryo-EM) has emerged as a transformative tool in structural biology, enabling visualization of macromolecular complexes at atomic or near-atomic resolution without crystallization (Kühlbrandt 2014; Nogales 2016; Renaud et al. 2018). Central to cryo-EM is the computational task of reconstructing a 3D volume from a large collection of 2D projection images, each corresponding to a different, unknown viewing direction of identical particles embedded in vitreous ice.
This inverse problem is inherently ill-posed and poses significant computational challenges. First, cryo-EM images are severely corrupted by noise due to the low electron dose required to prevent radiation damage. For experimental datasets, the signal-to-noise (SNR) could be as low as around 20-20 dB (Bendory, Bartesaghi, and Singer 2020; Bepler et al. 2020). Second, the orientations (poses) of individual particles are unknown and must be inferred jointly with the 3D structure. Third, many biological samples exhibit structural heterogeneity, with multiple conformational states coexisting in the same dataset.
@@ Introduction
Approaches to homogeneous reconstruction include methods based on backprojection, iterative expectation-maximization with voxel-based volumes (Tang et al. 2007; Scheres 2012; Punjani et al. 2017; Shekarforoush et al. 2024), and more recently, neural representation learning (Zhong et al. 2021a, b), which models the 3D volume using coordinate-based networks. In parallel, Gaussian mixture models (GMMs) have received attention for their continuous, compact, and physically interpretable parameterization of molecular density (Chen and Ludtke 2021; Chen, Schmid, and Chiu 2023). Notably, GMMs offer a natural connection to atomic models and can represent fine structural details using fewer parameters (Chen, Toader, and Roy 2023; Li et al. 2024; Schwab et al. 2024; Chen 2025).
Despite their conceptual appeal, existing GMM-based methods for cryo-EM reconstruction rely on nontrivial prerequisite steps. These approaches (2021; 2023; 2023; 2024; 2024; 2025) typically rely on consensus volumes from external pipelines, or even atomic models, for initialization, and have not demonstrated stable convergence when directly optimizing from experimental images. In fact, no prior method achieves reliable GMM-based reconstruction even under known particle poses, due to the inherent difficulty of optimizing mixture parameters in extreme noise. As a result, GMMs lack a self-contained and stable formulation that can serve as a backbone for broader reconstruction workflows.
In this work, we propose cryoSplat, a GMM-based homogeneous reconstruction method that fills this foundational gap. Given known particle poses, cryoSplat performs stable and efficient reconstruction directly from raw cryo-EM projection images, starting from random initialization and requiring no external priors. Inspired by recent advances in 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023), we model the 3D density as a mixture of anisotropic Gaussians and project them into 2D using a novel differentiable orthographic splatting algorithm consistent with cryo-EM physics. To support practical scalability and training efficiency, we develop a CUDA-accelerated real-space renderer that enables fast rasterization and optimization of the GMM.
@@
  • A self-contained GMM-based reconstruction method: We present cryoSplat as the first method capable of performing cryo-EM homogeneous reconstruction from scratch using only a Gaussian Mixture Model, thereby establishing the missing foundation needed to develop GMMs into standalone reconstruction tools.
@@
Related Work
Volume Representation in Cryo-EM
In cryo-EM experiments, purified biomolecules are rapidly frozen in a thin layer of vitreous ice, where each particle adopts a random orientation. A high-energy electron beam passes through the specimen, interacts with the electrostatic potential of the particles, and is recorded on a detector as a 2D projection image (Singer and Sigworth 2020). The goal of cryo-EM reconstruction is to recover the 3D electrostatic potential, i.e., the volume, from a large set of such noisy and randomly oriented 2D projections. Central to cryo-EM reconstruction is the choice of volume representation.
Voxel-based representations
are the most widely used in conventional cryo-EM software, e.g., RELION (2016), cryoSPARC (2017) and EMAN2 (2007). The 3D volume is discretized into a regular grid of density values, enabling fast projection and reconstruction via FFT-based algorithms. Despite their practical success, voxel grids are memory-intensive and inherently discrete, which limits their compatibility with modern learning-based analysis frameworks.
Neural fields
represent the volume as a continuous function parameterized by neural networks. These methods (Zhong et al. 2021a, b; Levy et al. 2022a, b, 2025) offer differentiability, implicit smoothness, and natural compatibility with learning-based heterogeneous analysis. However, the implicit nature of neural fields often comes at the cost of interpretability, and such models are typically slow to train and difficult to constrain with biological priors.
Gaussian mixture models
have a long history in structural biology, with early uses for molecular approximation (Grant and Pickup 1995; Grant, Gallardo, and Pickup 1996; Kawabata 2008). E2GMM (Chen and Ludtke 2021) was among the first to apply GMMs to cryo-EM heterogeneous reconstruction. Like neural fields, GMMs can approximate any smooth density function and support differentiable optimization. More importantly, GMMs provide an explicit and interpretable representation that naturally links to atomic structures. Recent studies (Chen, Schmid, and Chiu 2023; Chen, Toader, and Roy 2023; Li et al. 2024; Schwab et al. 2024; Chen 2025) have shown that GMMs can capture molecular flexibility by modeling atomic motion directly, making them highly suitable for heterogeneous reconstruction and downstream structural analysis.
@@
Gaussian Splatting
3DGS (Kerbl et al. 2023) is a recent differentiable rendering technique developed for real-time novel view synthesis. It represents a 3D scene as a collection of anisotropic Gaussians and renders images via rasterization-based accumulation and alpha blending (Zwicker et al. 2002). While 3DGS achieves high visual fidelity in synthetic and real-world RGB datasets, as a volume rendering method, it is not a physically accurate model of natural image formation (Huang et al. 2024).
@@
Building upon recent progress in differentiable volume rendering, particularly the Gaussian splatting framework introduced in 3DGS (2023), we adopt a tile-based rasterization strategy to enable scalable and efficient computation. However, the original 3DGS formulation is not directly applicable to cryo-EM due to several fundamental mismatches: (i) it employs perspective projection consistent with a pinhole camera model, in contrast to the orthographic projection in cryo-EM imaging; (ii) it prioritizes photorealistic novel view synthesis over physical reconstruction accuracy; and (iii) its coordinate system is misaligned with the Fourier slice theorem, which underpins cryo-EM reconstruction.
@@
To address these issues, cryoSplat introduces several key adaptations: (i) we replace the heuristic alpha compositing with physically grounded line integrals to reflect the transmission nature of electron imaging; (ii) we fix the normalization between 3D-to-2D transformation and apply consistent learning rates across all parameters to ensure stable optimization; and (iii) we align the rasterization coordinate system with the FFT grid, allowing accurate gradient propagation through CTF modulation.
@@
These modifications collectively enable cryoSplat to perform stable, end-to-end differentiable reconstruction from raw cryo-EM particle images, starting from random initialization without relying on externally provided volumes.
Image Formation
@@
As shown in Fig. 1(b), a beam of electrons passes through a vitrified specimen, and the transmitted wavefront undergoes phase shifts due to the specimen’s electrostatic potential(Singer and Sigworth 2020). Under the weak phase approximation, these phase shifts are linearly related to the 3D potential (volume), and the image formed at the detector is a line integral (projection) of this potential along the beam direction, further convolved with $\mathcal{H}$, the point spread function (PSF) of the imaging system.
@@
In homogeneous reconstruction, it is assumed that all particle images $\mathcal{I}$ correspond to identical copies of a single 3D volume $\mathcal{V}:\mathbb{R}^{3}\rightarrow\mathbb{R}$, and that any conformational or compositional heterogeneity is negligible. Under this assumption, the image formation model can be expressed as:
$\mathcal{I}(r_{x},r_{y})=\mathcal{H}(r_{x},r_{y})*\int_{\mathbb{R}}\mathcal{V}(\mathbf{W}^{\top}\mathbf{r}+\mathbf{t})dr_{z}+\epsilon,$ (1)
@@
where $\mathbf{r}=[r_{x},r_{y},r_{z}]^{\top}$ denotes the 3D Cartesian coordinates in real space, $\mathbf{W}\in SO(3)$ is the 3D pose of the volume, and $\mathbf{t}=[t_{x},t_{y},0]^{\top}$ is the in-plane translation, accounting for imperfect centering during particle cropping. The noise term $\epsilon$ is typically modeled as independent, zero-centered Gaussian noise.
According to the Fourier slice theorem (Bracewell 1956), illustrated in Fig. 1(c), the 2D Fourier transform of a projection corresponds to a central slice of the 3D Fourier transform of the volume, orthogonal to the projection direction and passing through the origin. Based on this property, an alternative and widely adopted formulation models reconstruction directly in the Fourier domain, where the image formation model becomes:
$\widehat{\mathcal{I}}(k_{x},k_{y})=\widehat{\mathcal{H}}(k_{x},k_{y})\cdot\widehat{\mathcal{V}}(\mathbf{W}^{\top}\mathbf{k})\cdot e^{-2\pi i\mathbf{k}^{\top}\mathbf{t}}+\widehat{\epsilon}$, (2)
where $\mathbf{k}=[k_{x},k_{y},0]^{\top}$ denotes the Cartesian coordinates in Fourier space, and the 2D spectrum $\widehat{\mathcal{I}}$, the CTF $\widehat{\mathcal{H}}$ and the 3D spectrum $\widehat{\mathcal{V}}$ denote the Fourier transform of $\mathcal{I}$, $\mathcal{H}$ and $\mathcal{V}$, respectively. The noise term $\widehat{\epsilon}$ is similarly modeled as independent, zero-mean Gaussian noise in the Fourier domain.
CryoSplat
Anisotropic GMMs
@@
are developed to represent the volume, which can be written in the form
$\mathcal{V}(\mathbf{r})=\sum_{i=1}^{N}A_{i}\mathcal{G}_{i}(\mathbf{r}),$ (3)
where $N$ denotes the number of Gaussians and $A_{i}$ is the amplitude of the $i$-th normalized Gaussian $\mathcal{G}_{i}(\mathbf{r})$.
@@
$\mathcal{I}(r_{x},r_{y})=\mathcal{H}(r_{x},r_{y})*\sum_{i=1}^{N}A_{i}\int_{\mathbb{R}}\mathcal{G}_{i}(\mathbf{W}^{\top}\mathbf{r}+\mathbf{t})dr_{z}.$ (4)
@@
$\mathcal{G}(\mathbf{r}|\boldsymbol{\mu},\boldsymbol{\Sigma})=\frac{1}{(2\pi)^{\frac{3}{2}}|\boldsymbol{\Sigma}|^{\frac{1}{2}}}\exp(-\frac{1}{2}(\mathbf{r}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}^{-1}(\mathbf{r}-\boldsymbol{\mu})),$ (5)
where $\boldsymbol{\mu}\in\mathbb{R}^{3}$ and $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$ denote the mean (center) and the covariance matrix, respectively. The determinant $|\boldsymbol{\Sigma}|$ ensures proper normalization. Following (Kerbl et al. 2023), to guarantee the positive semidefinite property, we construct the covariance matrix as:
$\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top},$ (6)
where $\mathbf{S}=\mathrm{diag}(\mathbf{s})$ is a diagonal scaling matrix and $\mathbf{R}\in SO(3)$ is a rotation matrix. In our implementation, we store the scaling vector $\mathbf{s}=[s_{x},s_{y},s_{z}]^{\top}$ and parameterize $\mathbf{R}$ using a quaternion $\mathbf{q}=[q_{w},q_{x},q_{y},q_{z}]^{\top}$. To ensure positivity and stable gradients during optimization, we apply a softplus function to both the amplitude $A$ and the scaling vector $\mathbf{s}$. The quaternion $\mathbf{q}$ is normalized to ensure it represents a valid rotation. Altogether, each anisotropic Gaussian is parameterized by the 11-dimensional set $\{\mu_{x},\mu_{y},\mu_{z},s_{x},s_{y},s_{z},q_{w},q_{x},q_{y},q_{z},A\}$.
When the scaling is isotropic, i.e., $s_{x}=s_{y}=s_{z}=\sigma$, the anisotropic Gaussian reduces to the standard isotropic form:
$\mathcal{G}(\mathbf{r}|\boldsymbol{\mu},\sigma)=\frac{1}{(2\pi)^{\frac{3}{2}}\sigma}\exp(-\frac{\|\mathbf{r}-\boldsymbol{\mu}\|_{2}^{2}}{2\sigma^{2}}),$ (7)
which is more commonly used in previous cryo-EM reconstruction works (2021; 2023; 2023; 2024; 2024; 2025).
The viewing transformation
is the first step in simulating image formation, aligning each Gaussian with a given projection direction. Since the parameters $\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}$ describe Gaussians in world coordinates, we must transform them into the image-space coordinates before projection.
According to the derivation in (Zwicker et al. 2002), applying an affine transformation to a Gaussian results in another Gaussian with appropriately transformed parameters. In our case, the transformation consists of a rotation $\mathbf{W}\in SO(3)$ and a 2D in-plane translation $\mathbf{t}\in\mathbb{R}^{3}$, leading to:
$\dot{\mathcal{G}}(\mathbf{r}|\dot{\boldsymbol{\mu}},\dot{\boldsymbol{\Sigma}})=\mathcal{G}(\mathbf{W}^{\top}\mathbf{r}+\mathbf{t}|\boldsymbol{\mu},\boldsymbol{\Sigma}),$ (8)
where the transformed mean and covariance are given by $\dot{\boldsymbol{\mu}}=\mathbf{W}\boldsymbol{\mu}+\mathbf{t}$ and $\dot{\boldsymbol{\Sigma}}=\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}$, respectively.
The orthogonal projection

### v3

@@ Abstract
As a critical modality for structural biology, cryogenic electron microscopy (cryo-EM) facilitates the determination of macromolecular structures at near-atomic resolution. The core computational task in single-particle cryo-EM is to reconstruct the 3D electrostatic potential of a molecule from noisy 2D projections acquired at unknown orientations. Gaussian mixture models (GMMs) provide a continuous, compact, and physically interpretable representation for molecular density and have recently gained interest in cryo-EM reconstruction. However, existing methods rely on external consensus maps or atomic models for initialization, limiting their use in self-contained pipelines. In parallel, differentiable rendering techniques such as Gaussian splatting have demonstrated remarkable scalability and efficiency for volumetric representations, suggesting a natural fit for GMM-based cryo-EM reconstruction. However, off-the-shelf Gaussian splatting methods are designed for photorealistic view synthesis and remain incompatible with cryo-EM due to mismatches in the image formation physics, reconstruction objectives, and coordinate systems. Addressing these issues, we propose cryoSplat, a GMM-based method that integrates Gaussian splatting with the physics of cryo-EM image formation. In particular, we develop an orthogonal projection-aware Gaussian splatting, with adaptations such as a view-dependent normalization term and FFT-aligned coordinate system tailored for cryo-EM imaging. These innovations enable stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images using random initialization. Experimental results on real datasets validate the effectiveness and robustness of cryoSplat over representative baselines. The code will be released upon publication.
1 Introduction
Single particle cryogenic electron microscopy (cryo-EM) has emerged as a transformative tool in structural biology, enabling visualization of macromolecular complexes at atomic or near-atomic resolution without crystallization (Kühlbrandt, 2014; Nogales, 2016; Renaud et al., 2018). Central to cryo-EM is the computational task of reconstructing a 3D volume from a large collection of 2D projection images, each corresponding to a different, unknown viewing direction of identical particles embedded in vitreous ice.
This inverse problem is inherently ill-posed and computationally challenging. First, cryo-EM images are severely corrupted by noise due to the low electron dose required to prevent radiation damage. For experimental datasets, the signal-to-noise (SNR) could be as low as around 20-20 dB (Bendory et al., 2020; Bepler et al., 2020). Second, the orientations (poses) of individual particles are unknown and must be inferred jointly with the 3D structure. Third, many biological samples exhibit structural heterogeneity, with multiple conformational states coexisting in the same dataset.
@@ 1 Introduction
Approaches to homogeneous reconstruction include methods based on backprojection, iterative expectation-maximization with voxel-based volumes (Tang et al., 2007; Scheres, 2012; Punjani et al., 2017; Shekarforoush et al., 2024), and more recently, neural representation learning (Zhong et al., 2021a; b), which models the 3D volume using coordinate-based networks. In parallel, Gaussian mixture models (GMMs) have received attention for their continuous, compact, and physically interpretable parameterization of molecular density (Chen & Ludtke, 2021; Chen et al., 2023a). Notably, GMMs offer a natural connection to atomic models and can represent fine structural details using fewer parameters (Chen et al., 2023b; Li et al., 2024; Schwab et al., 2024; Chen, 2025).
Despite their conceptual appeal, existing GMM-based methods (Chen & Ludtke, 2021; Chen et al., 2023a; b; Li et al., 2024; Schwab et al., 2024; Chen, 2025) for cryo-EM reconstruction rely on nontrivial prerequisite steps. These approaches typically rely on consensus volumes from external pipelines, or even atomic models, for initialization, and have not demonstrated stable convergence when directly optimizing from experimental images. In fact, no prior method achieves reliable GMM-based reconstruction even under known particle poses, due to the inherent difficulty of optimizing mixture parameters in extreme noise. As a result, GMMs lack a self-contained and stable formulation that can serve as a backbone for broader reconstruction workflows.
In this work, we propose cryoSplat, a GMM-based homogeneous reconstruction method that fills this foundational gap. Given known particle poses, cryoSplat performs stable and efficient reconstruction directly from raw cryo-EM projection images, starting from random initialization and requiring no external priors. Inspired by recent advances in 3D Gaussian Splatting (3DGS) by Kerbl et al. (2023), we model the 3D density as a mixture of anisotropic Gaussians and project them into 2D using a novel differentiable orthographic splatting algorithm consistent with cryo-EM physics. To support practical scalability and training efficiency, we develop a CUDA-accelerated real-space renderer that enables fast rasterization and optimization of the GMM.
@@
  • A self-contained GMM-based reconstruction method: We present cryoSplat as the first method capable of performing cryo-EM homogeneous reconstruction from a randomly initialized Gaussian mixture model without an external prior, thereby establishing the missing foundation needed to develop GMMs into standalone reconstruction tools.
@@
2 Related work
2.1 Volume representation in cryo-EM
In cryo-EM experiments, purified biomolecules are rapidly frozen in a thin layer of vitreous ice, where each particle adopts a random orientation. A high-energy electron beam passes through the specimen, interacts with the electrostatic potential of the particles, and is recorded on a detector as a 2D projection image (Singer & Sigworth, 2020). The goal of cryo-EM reconstruction is to recover the 3D electrostatic potential, i.e., the volume, from a large set of such noisy and randomly oriented 2D projections. Central to cryo-EM reconstruction is the choice of volume representation.
2.1.1 Voxel-based representation
Voxel-based representations are the most widely used in conventional cryo-EM software, e.g., RELION (Scheres, 2016b), cryoSPARC (Punjani et al., 2017) and EMAN2 (Tang et al., 2007). The 3D volume is discretized into a regular grid of density values, enabling fast projection and reconstruction via FFT-based algorithms. Despite their practical success, voxel grids are memory-intensive and inherently discrete, which limits their compatibility with modern learning-based analysis frameworks.
2.1.2 Neural field
Neural fields represent the volume as a continuous function parameterized by neural networks. These methods (Zhong et al., 2021a; b; Levy et al., 2022a; b; 2025) offer differentiability, implicit smoothness, and natural compatibility with learning-based heterogeneous analysis. However, the implicit nature of neural fields often comes at the cost of interpretability, and such models are typically slow to train and difficult to constrain with biological priors.
2.1.3 Gaussian mixture model
Gaussian mixture models have a long history in structural biology, with early uses for molecular approximation (Grant & Pickup, 1995; Grant et al., 1996; Kawabata, 2008). E2GMM (Chen & Ludtke, 2021) was among the first to apply GMMs to cryo-EM heterogeneous reconstruction. Like neural fields, GMMs can approximate any smooth density function and support differentiable optimization. More importantly, GMMs provide an explicit and interpretable representation that naturally links to atomic structures. Recent studies (Chen et al., 2023a; b; Li et al., 2024; Schwab et al., 2024; Chen, 2025) have shown that GMMs can capture molecular flexibility by modeling atomic motion directly, making them highly suitable for heterogeneous reconstruction and downstream structural analysis.
@@ 2.1.3 Gaussian mixture model
2.2 Gaussian splatting
3DGS (Kerbl et al., 2023) is a recent differentiable rendering technique developed for real-time novel view synthesis. It represents a 3D scene as a collection of anisotropic Gaussians and renders images via rasterization-based accumulation and alpha blending (Zwicker et al., 2002). While 3DGS achieves high visual fidelity in synthetic and real-world RGB datasets, as a volume rendering method, it is not a physically accurate model of natural image formation (Huang et al., 2024).
@@ 3.1 Overview
Building upon recent progress in differentiable volume rendering, particularly the Gaussian splatting framework by Kerbl et al. (2023), we adopt a tile-based rasterization strategy for scalable and efficient computation. However, the original 3DGS formulation is not directly applicable to cryo-EM due to several fundamental mismatches: (i) it employs perspective projection consistent with a pinhole camera model, in contrast to the orthographic projection in cryo-EM imaging; (ii) it prioritizes photorealistic novel view synthesis over physical reconstruction accuracy; and (iii) its coordinate system is misaligned with the Fourier slice theorem, which underpins cryo-EM reconstruction.
@@
To address these issues, cryoSplat introduces several key adaptations: (i) we replace heuristic alpha blending with physically grounded line integrals to reflect the transmission nature of electron imaging; (ii) we fix the normalization between 3D-to-2D transformation and apply consistent learning rates across all parameters to ensure stable optimization; and (iii) we align the rasterization coordinate system with the FFT grid, allowing accurate gradient propagation through CTF modulation.
@@
These modifications collectively enable cryoSplat to perform stable, end-to-end differentiable reconstruction from raw cryo-EM particle images, starting from random initialization without relying on externally provided volumes or atomic models.
3.2 Image formation
@@
As shown in Fig. 1(b), electrons traverse a vitrified specimen, and the transmitted wavefronts undergo phase shifts due to the specimen’s electrostatic potential (Singer & Sigworth, 2020). Under the weak phase approximation, the phase shifts are linearly related to the 3D potential (volume), and the image formed at the detector is a line integral (projection) of this potential along the beam direction, further convolved with $H\!:{\mathbb{R}}^{2}\!\rightarrow\!{\mathbb{R}}$, the point spread function (PSF) of the imaging system.
@@
In homogeneous reconstruction, it is assumed that all particle images $Y:{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ correspond to identical copies of a single 3D volume $V:{\mathbb{R}}^{3}\rightarrow{\mathbb{R}}$, and that any conformational or compositional heterogeneity is negligible. Under this assumption, the image formation model can be expressed as:
$Y(r_{x},r_{y})=H(r_{x},r_{y})*\int_{{\mathbb{R}}}V({\bm{W}}^{\top}{\bm{r}}+{\bm{t}})dr_{z}+{\textnormal{$\epsilon$}},$ (1)
@@
where ${\bm{r}}=[r_{x},r_{y},r_{z}]^{\top}$ are the 3D Cartesian coordinates in real space, ${\bm{W}}\in\mathrm{SO}(3)$ is the 3D pose of the particle, and ${\bm{t}}=[t_{x},t_{y},0]^{\top}$ is the in-plane translation, accounting for imperfect centering during particle cropping. The noise term $\epsilon$ is modeled as independent, zero-mean Gaussian noise.
3.3 CryoSplat
3.3.1 Anisotropic GMM
@@
Anisotropic GMMs are developed to represent the volume, which can be written in the form
$V({\bm{r}})=\sum_{i=1}^{N}A_{i}G_{i}({\bm{r}}),$ (2)
where $N$ denotes the Guassian count and $A_{i}$ is the amplitude of the $i$-th normalized Gaussian $G_{i}({\bm{r}})$.
@@
$X(r_{x},r_{y})=H(r_{x},r_{y})*\sum_{i=1}^{N}A_{i}\int_{{\mathbb{R}}}G_{i}({\bm{W}}^{\top}{\bm{r}}+{\bm{t}})dr_{z}.$ (3)
@@
$G({\bm{r}}|{\bm{\mu}},{\bm{\Sigma}})=\frac{1}{(2\pi)^{\frac{3}{2}}|{\bm{\Sigma}}|^{\frac{1}{2}}}\exp\Big(-\frac{1}{2}({\bm{r}}-{\bm{\mu}})^{\top}{\bm{\Sigma}}^{-1}({\bm{r}}-{\bm{\mu}})\Big),$ (4)
where ${\bm{\mu}}\in{\mathbb{R}}^{3}$ and ${\bm{\Sigma}}\in{\mathbb{R}}^{3\times 3}$ denote the mean (position) and the covariance matrix (shape), respectively. The determinant $|{\bm{\Sigma}}|$ ensures proper normalization. Following Kerbl et al. (2023), to guarantee the positive semidefinite property, we construct the covariance matrix as:
${\bm{\Sigma}}={\bm{R}}{\bm{S}}{\bm{S}}^{\top}{\bm{R}}^{\top},$ (5)
where ${\bm{S}}=\mathrm{diag}({\bm{s}})$ is a diagonal scaling matrix and ${\bm{R}}\in\mathrm{SO}(3)$ is a rotation matrix. In our implementation, we store the scaling vector ${\bm{s}}=[s_{x},s_{y},s_{z}]^{\top}$ and parameterize $\bm{R}$} using a quaternion ${\bm{q}}=[q_{w},q_{x},q_{y},q_{z}]^{\top}$. To ensure positivity and stable gradients during optimization, we apply a softplus function to both the amplitude $A$ and the scaling vector $\bm{s}$}. The quaternion $\bm{q}$} is normalized to ensure it represents a valid rotation. Altogether, each anisotropic Gaussian is parameterized by the 11-dimensional set $\{\mu_{x},\mu_{y},\mu_{z},s_{x},s_{y},s_{z},q_{w},q_{x},q_{y},q_{z},A\}$.
3.3.2 Viewing transformation
The viewing transformation is the first step in simulating image formation, aligning each Gaussian with a given projection direction. Since the parameters $\bm{\mu}$} and $\bm{\Sigma}$} describe Gaussians in world coordinates, we must transform them into the image-space coordinates before projection.
According to the derivation in Zwicker et al. (2002), applying an affine transformation to a Gaussian results in another Gaussian with appropriately transformed parameters. In our case, the transformation consists of a rotation ${\bm{W}}\in\mathrm{SO}(3)$ and a 2D in-plane translation ${\bm{t}}\in{\mathbb{R}}^{3}$, leading to:
$\dot{G}({\bm{r}}|\dot{{\bm{\mu}}},\dot{{\bm{\Sigma}}})=G({\bm{W}}^{\top}{\bm{r}}+{\bm{t}}|{\bm{\mu}},{\bm{\Sigma}}),$ (6)
where the transformed mean and covariance are given by $\dot{{\bm{\mu}}}={\bm{W}}({\bm{\mu}}-{\bm{t}})$ and $\dot{{\bm{\Sigma}}}={\bm{W}}{\bm{\Sigma}}{\bm{W}}^{\top}$.
3.3.3 Orthogonal projection